Five years ago, Karma was one third as old as her father was then. In five years, Karma will be half as old as her father will be then. Find their ages now.
Answer by lwsshak3(11628)

Five years ago, Karma was one third as old as her father was then. In five years, Karma will be half as old as her father will be then. Find their ages now.
***
5 yrs ago:
let x=Karma's age
3x=father's age
..
now:
x+5=Karma's age
3x+5=father's age
..
In 5 yrs:
x+5+5=x+10=Karma's age
3x+5+5=3x+10=father's age
3x+10=2(x+10) (In five years, Karma will be half as old as her father will be then.)
3x+10=2x+20
x=10
x+5=15
3x+5=35
..
their ages now:
Karma's age=15
father's age=35